OpenStudy (anonymous):

bio gas is a mixture of gases.according to its volume 85% C2H6, 6% CH4, 5% CO , 2% CO2 , 0.5% N2 and remaining is deactivated gas. Bio gas is burnt in a cylinder of 0.25m^3 volume. then the emitted heat is used to boil 100l of water in an aluminium vessel of 10kg.At the beginning temperature was 25C . At the end it was 100C in the vessel. 20% of the heat emitted was wasted. S.H.C of water is 4200 S.H.C of Al is 2666.67 C2H6 + 7/2 O2 =====2CO2 + 3H2O -4500kJmol CH4 + 2O2 ====CO2 + 2H2O -2000kJmol CO + 1/2 O2 ==== CO2 -1500kJmol 1. Find the emitted heat,No. of moles

OpenStudy (aaronq):

They gave you a lot of information thats not necessary to answer the question. q=80% of heat emitted \(\dfrac{q}{0.8}=\dfrac{q_T}{1}\) so \(q_T=\dfrac{q}{0.8}\) \(q_T=\dfrac{q}{0.8}=\underbrace{m*C_P*\Delta T}_{water}+\underbrace{m*C_P*\Delta T}_{Al}\)

OpenStudy (aaronq):

\(q_T=\dfrac{\underbrace{m*C_p*\Delta T}_{water}+\underbrace{m*C_p*\Delta T}_{Al}}{0.8}\) that last equation should've read. and they want the number of moles of what?

OpenStudy (anonymous):

Actually I don't know what no. of moles they want? It says Total moles I think it's total moles in bio gas mixture When calculating heat emitted why is it divided by 0.8? They ask total heat that emitted though 20% is wasted it's also heat emitted right? Please reply!

OpenStudy (aaronq):

because we want to know 100% of the heat emitted, by dividing by 0.8 you find that. (look at the first ratio i posted). hm it seems that you could just take the weighted average of the heat evolved, so: q= [85%*(-4500 kJ/mol)+6%*(-2000 kJ/mol)+ 5%(-1500 kJ/mol)]*n

OpenStudy (anonymous):

Finding the no. of moles is the problem for me

OpenStudy (anonymous):

Also this is a part of a question the details given are needed for the other parts

OpenStudy (aaronq):

unless the percentages are given in mass, which then you'd have to assume some mass to correct for that.

OpenStudy (anonymous):

i thought of that but N2 too is burnt right? they have not given a reaction for that but a certain amount of energy is emitted through it also right?

OpenStudy (aaronq):

\(N_2\) will likely stay as \(N_2\)

OpenStudy (aaronq):

it's very inert

OpenStudy (anonymous):

in the next part they ask no of moles of CO,CO2,N2,CH4,C2H6 and deactivated gas separately in the cylinder.How do you find that?

OpenStudy (aaronq):

i think you might need to do some more other stuff to account for moles, because whats given might be in mass.

OpenStudy (anonymous):

i'm pretty sure that its given in volume can't we say that volume is directly proportional to moles( Avogadro law)

OpenStudy (aaronq):

OH it is definitely in volume. Then yes, proceed with: q= [85%*(-4500 kJ/mol)+6%*(-2000 kJ/mol)+ 5%(-1500 kJ/mol)]*n

OpenStudy (aaronq):

then you wanna do another ratio to find the total number of moles n/96%=\(n_T\)/100%

OpenStudy (anonymous):

I didn't understand it I found the total no. of moles = 8.33 mol to find moles of each gas can we multiply it by there volume percentages?

OpenStudy (aaronq):

yes